(TCO D) XYZ company paid a dividend of $4.37 in the past 12 months, and it has a growth rate of 6.5 percent. The stock currently sells for $175.

Using the constant-growth DDM, calculate the current price of the stock. You must show your work.
Answer #1

Given that the recent dividend is $4.37
Growth rate is 6.5%
Expected dividend is (Recent dividend)*(1+Growth rate)=4.37*(1+6.5%)=4.65405
Current price of the stock=$175

According to the constant growth DDM model:
Current price = (Expected Dividend)/(Required Return - growth rate)
Current price and growth rate are already given, so we can determine the required return using the constant growth dividend discount model formula
Now, current price = (Expected Dividend)/(Required Return - growth rate)
=>175= (4.65405)/(Required Return - 6.5%)
=>(Required Return - 6.5%)=(4.65405)/175
=>Required Return =(4.65405)/175 +6.5%
=>Required Return =0.026594571 +6.5%
=0.091594571 or 9.16% (Rounded to two decimal places)